OpenStudy (anonymous):

Anybody want to check my answers for a medal,fan and me loving you forever?!

OpenStudy (anonymous):

Given the functions m(x) = 4x − 11 and n(x) = x − 10, solve m[n(x)] and select the correct answer below. A. m[n(x)] = 4x − 51 B. m[n(x)] = 4x − 29 C. m[n(x)] = 4x^2− 51 D. m[n(x)] = 4x^2 − 29 I got b

OpenStudy (anonymous):

Given the functions f(x) = 6x + 11 and g(x) = x + 6, which of the following functions represents f[g(x)] correctly? f[g(x)] = 6x + 47 f[g(x)] = 6x + 17 f[g(x)] = 6x2 + 47 f[g(x)] = 6x2 + 17 I got A!

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle m[\color{red}{n(x)}] =4(\color{red}{x-10})-11 }\) first question is NOT b.

OpenStudy (anonymous):

Yea i did that and i got 4x-40-11... and -40-11 is -29 so my final answer was: 4x-29

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f[\color{red}{g(x)}] =6(\color{red}{x+6})+11 }\) 2nd question is done correctly, nice job!

OpenStudy (solomonzelman):

And no, sorry, but the first question is still incorrect

OpenStudy (anonymous):

Would it be d then cuz its for sure 29 im just not sure if there is an exponent?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle m[\color{red}{n(x)}] =4(\color{red}{x-10})-11 }\) \(\large\color{black}{ \displaystyle m[\color{red}{n(x)}] =4x-40-11 }\)

OpenStudy (solomonzelman):

Let me ask you, what is 40+11=?

OpenStudy (anonymous):

51 but isn't it 40-11?

OpenStudy (solomonzelman):

yes, it is -51:)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle m[\color{red}{n(x)}] =4x-51 }\)

OpenStudy (anonymous):

but isn't it 40-11, which is 29?

OpenStudy (anonymous):

you said m[n(x)]=4x−40−11... so 40-11 is 29 right? Like why did you add cuz now im confused.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle m[\color{red}{n(x)}] =4x-40-11 }\) -40-11 = -51 No?

OpenStudy (anonymous):

ohhh I see!! Ok cuz i was just doing 40-11 instead of -40 -11

OpenStudy (solomonzelman):

yeah, that is what you were doing, but now you got it i see(?) any questuions regarding any of the two problems you posted?

OpenStudy (anonymous):

Thank you so much!! No, can you check my answers for 3 more?

OpenStudy (solomonzelman):

Yes, but in separate questions,, please. cu

OpenStudy (anonymous):

ok you mean one at a time?

OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve (f + g)(3) and select the correct answer below: f(x) = 2x + 21 g(x) = x − 24 A. −27 B. −21 C. 6------- my answer! D. 48

OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve f[g(10)] and select the correct answer below: f(x) = 10x + 8 g(x) = x + 9 A. 2,052 B. 98 C. 190 D. 198--------my answer!

OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve (f ⋅ g)(3) and select the correct answer below: f(x) = 4x2 + 12 g(x) = x − 1 A. 47 B.−47 C. 96------my answer D. −96

OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve (f/g) (−5) and select the correct answer below: f(x) = 2x − 20 g(x) = x − 1 A. −5 B. 5-------my answer! C. 1/6 D. 30

OpenStudy (anonymous):

I think the rest are right

OpenStudy (solomonzelman):

#3 is correct

OpenStudy (anonymous):

ok thank you!

OpenStudy (solomonzelman):

#4 is correct

OpenStudy (solomonzelman):

#5 is correct as well. Nice work!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (solomonzelman):

Sure, ....